# Tracking-Assisted Object Detection with Event Cameras

### v1

@@ Abstract
Event-based object detection has recently garnered attention in the computer vision community due to the exceptional properties of event cameras, such as high dynamic range and no motion blur. However, feature asynchronism and sparsity cause invisible objects due to no relative motion to the camera [^1] The objects with no relative motion to the camera, still objects, invisible objects, and pseudo-occluded objects are used interchangeably in the paper., posing a significant challenge in the task. Prior works have studied various memory mechanisms to preserve as many features as possible at the current time, guided by temporal clues. While these implicit-learned memories retain some short-term information, they still struggle to preserve long-term features effectively. In this paper, we consider those invisible objects as pseudo-occluded objects and aim to reveal their features. Firstly, we introduce visibility attribute of objects and contribute an auto-labeling algorithm to append additional visibility labels on an existing event camera dataset. Secondly, we exploit tracking strategies for pseudo-occluded objects to maintain their permanence and retain their bounding boxes, even when features have not been available for a very long time. These strategies can be treated as an explicit-learned memory guided by the tracking objective to record the displacements of objects across frames. Lastly, we propose a spatio-temporal feature aggregation module to enrich the latent features and a consistency loss to increase the robustness of the overall pipeline. We conduct comprehensive experiments to verify our method’s effectiveness where still objects are retained but real occluded objects are discarded. The results demonstrate that (1) the additional visibility labels can assist in supervised training, and (2) our method outperforms state-of-the-art approaches with a significant improvement of 7.9% absolute mAP.
@@ 1 Introduction
Event cameras measure only per-pixel intensity changes asynchronously and are suitable for detecting fast-moving objects. For those slow-moving objects or still objects, it is very hard to notice their physical existence if there are no temporal clues. For example, Fig. 1 shows a video sequence from 1 Megapixel Automotive Detection Dataset where two cars slow down from Fig. 1(a) to Fig. 1(b), remain still from Fig. 1(c) to Fig. 1(d), and start moving at Fig. 1(e). The features in the ground truth bounding boxes disappear gradually when slowing down, become empty when being still, and reappear gradually when speeding up. Human beings are capable of knowing the physical existence of these two cars as long as they observe the whole video sequence, even if there are no features in the ground truth bounding boxes. This kind of ability is known as object permanence[3, 2, 10]. In fact, object permanence is a term from developmental psychology that describes how infants know the continuous existence of unseen objects.
@@
Prior works propose different kinds of memory architectures to retain object permanence in event-based object detection. RED[34] and RVT[19] propose multi-stage convolution and recurrent vision transformer with different spatial resolutions, respectively, along with multi-stage ConvLSTMs. ASTMNet[26] proposes a temporal attention convolutional network (TACN) along with a single-stage ConvLSTM. DMANet[45] and HMNet[21] propose memories with different temporal resolutions to handle objects with different speeds, respectively. One common characteristic among these models is that they all use implicit memories with different spatial or temporal resolutions. After some experiments [^2] We cannot experiment on ASTMNet[26] due to no available source code. HMNet[21] is trained on GEN1 Automotive Detection Dataset with a smaller spatial resolution (304 x 240). Therefore, we downsample the spatial resolution of 1 Megapixel Automotive Detection Dataset (1280 x 720) and test them on HMNet[21] directly. shown in Fig. 1, we found that only RVT[19] can better handle the long-range dependency problem, i.e. when objects slow down and there are eventually no relative motions between objects and the camera, these objects become invisible regardless of their physical existence. Meanwhile, RVT[19] produces many false positives with high confidence scores due to its implicit ConvLSTMs. Other models in Fig. 1 are memoryless as long as the duration of missing features is too long, which is over 30 seconds from Fig. 1(c) to Fig. 1(d) and degrades the mAP performance of object detection.
@@
In order to supervise the training of object permanence, we need additional visibility labels to distinguish between moving objects and still objects. Hence, we propose an algorithm to perform the annotation automatically. Furthermore, our model is based on some joint object detection and tracking models, which give object detectors tracking ability. We augment these models with a spatio-temporal feature aggregation module and a consistency loss to not only enrich the latent features as shown in Fig. 1(g) but also increase the robustness of the overall pipeline where still objects are retained but real occluded objects are discarded. In summary, our model outperforms state-of-the-art event-based object detectors by 7.9% absolute mAP and our main contributions to this work are as follows:
@@
    We propose to make invisible objects visible by considering those invisible objects as pseudo-occlusion and exploiting tracking as an explicit memory to retain their bounding boxes despite having no feature for a very long time.
@@
    We propose an auto-labeling algorithm to annotate still objects, which are used to supervise the training of object permanence.
@@ 3 Method
Figure 2: The architecture of the proposed TEDNet.
@@ 3.1.3 Continuity Maintenance
is used to make the variation of visibility label for each specific tracking identity as smooth as possible and discard those bounding boxes with no features at their first occurrence. If there is no corresponding bounding box with the same tracking identity in the previous frame, a visibility label is determined by only the occupancy rate from Algorithm 1 to 1. Also, a buffer dubbed still_hits is needed to memorize the number of still hits for each tracking identity individually in the history and smooth out the impulse of visibility variation owing to the original label noises. If there is a corresponding tracked object in the previous frame, a visibility label is determined by either both occupancy rate and displacement from Algorithm 1 to 1 or only the still_hits buffer from Algorithm 1 to 1. Finally, not only visible objects but also those invisible objects that are tracked for two previous frames are added to the new labels from Algorithm 1 to 1. As a result, only $N_{i}$ out of $M_{i}$ objects are preserved with additional visibility labels for $i$-th frame.
@@
Input: $inputFeatures$ = tensor of shape $[T,\ C,\ H,\ W]$
         $oldLabels[i]=\{[x_{j},\ y_{j},\ w_{j},\ h_{j},\ trackId_{j}]\ |\ j\in[1,\ M_{i}]\},$
         $i\in[1,\ T]$
@@
Output:
         $newLabels[i]=\{[x_{k},\ y_{k},\ w_{k},\ h_{k},\ trackId_{k},\ visibility_{k}]\ |$
         $k\in[1,\ N_{i}]\},\ i\in[1,\ T]$
@@
5 $still\_hits\leftarrow\{\}$;
/* key-value pairs where values are limited to 5. */
6 $track\_vis\leftarrow\{\}$;
/* key-value pairs where values are queues with size 2. */
@@ 28                   $still\_hits[trackId_{j}]\leftarrow still\_hits[trackId_{j}]+1$;
31            if $visibility_{j}=1.0$ or ($track\_vis[trackId_{j}][0]=i-2$ and $track\_vis[trackId_{j}][1]=i-1$) then
@@ 3.3 Spatio-Temporal Feature Aggregation
The most significant challenge in event cameras is the long-range dependency problem. In this work, we propose to apply either a simple 3D convolutional network(C3D)[44] or a deformable 3D convolutional network(D3D)[49] as well as a ConvGRU in joint object detection and tracking module as our spatio-temporal feature aggregator. Unlike [44, 49] that input a video sequence to get a better spatio-temporal feature for the downstream task, our approach inputs only two frames, i.e., a current frame $I_{t}$ and a previous history frame $H_{t-1}$. The current history frame is obtained by $H_{t}=\emph{X3D}(I_{t},\ H_{t-1})$ where X3D is either C3D or D3D. To be more specific, we integrate a 3D convolution with a recurrent neural network to get a better spatio-temporal feature as an input to the following joint object detection and tracking module.
@@ 3.5 Consistency Loss
$min\Big{(}L_{con}\Big{)}=min\bigg{(}\Big{|}g(D^{t})\times h(C^{t})-V^{t}\Big{|}\bigg{)}=min\bigg{(}\Big{|}{\|D^{t}\|}_{2}\times e^{-relu(C^{t})}-V^{t}\Big{|}\bigg{)}$ (5)
$L_{con}=\frac{1}{N}\sum_{i=1}^{N}\bigg{|}{\Big{\|}D^{t}_{p_{i}}\Big{\|}}_{2}\times e^{-relu(C^{t}_{p_{i}})}-V^{t}_{p_{i}}\bigg{|}$ (6)
@@ 4.1 Dataset with Auto-Labeling
Prophesee, a company developing event-based sensors and algorithms, released the first large-scale and high-resolution (1280 $\times$ 720) event-based object detection dataset and its deep neural network RED[34] for automotive scenarios. This 1 Megapixel Automotive Detection Dataset is captured by the 1-megapixel event camera[17] and a standard RGB camera, but only event data and its annotations are released. The annotations come from the inference result of the RGB image and transfer to the event camera coordinate by geometric transformation. After visualizing the annotations in the dataset, we found that there are several incorrect bounding boxes with either imprecise locations and sizes or incorrect labels. Most importantly, several bounding boxes contain no features at the very beginning of some videos because they are still and visible by an RGB camera but invisible by an event camera, and humans cannot recognize their physical existence with only event data even though we have temporal clues. These incorrect bounding boxes limit the object detection performance and make model training awkward. Therefore, we clean the dataset by an auto-labeling algorithm in Sec. 3.1 with hyperparameters $D_{value}$=0.03 and $O_{value}$=0.1 to not only insert additional visibility labels to supervise the training of object permanence but also discard those bounding boxes that are impossible to be detected by humans. We named the dataset before/after data cleaning noisy GT/clean GT and its corresponding model noisy model/clean model.
@@ 4.2 Ablation Study
The models in ablation studies are trained on clean GT and tested on both noisy GT and clean GT as shown in Tab. 2. We choose two joint object detection and tracking models as our baselines, CenterTrack and PermaTrack[43], respectively. The CenterTrack shown here is the original one[50] with ConvGRU, making it comparable with PermaTrack[43].
4.2.1 Spatio-Temporal Feature Aggregation
4.2.2 Consistency Loss
We conduct consistency loss between displacement map $D^{t}$ and visibility map $V^{t}$ to model their proportional relationship. Consistency loss can only be applied to PermaTrack[43] because there is no visibility head in CenterTrack[50]. Comparing (d) with (g), the consistency loss enhances the mAP@0.5 by 2.3% on noisy GT and 2.5% on clean GT. Comparing (f) with (h), the consistency loss degrades the mAP@0.5 by 0.9% on noisy GT and 0.4% on clean GT if we use C3D as the spatio-temporal feature aggregation module. In contrast, comparing (e) with our TEDNet, the consistency loss improves the mAP@0.5 by 2.4% on noisy GT and 2.3% on clean GT if we use D3D as the spatio-temporal feature aggregation module. To sum up, consistency loss offers a better regularization to harness the advantages of deformable convolutions and increases the robustness of the overall pipeline.
@@ 4.2.2 Consistency Loss
* The CenterTrack here is the original one with ConvGRU.
@@ 4.4 Visualization of Detection Results
We visualize the detection results in Fig. 3 by a video with two cars remaining still for over 30 seconds and some other cars passing in front of them in between. For RED[34], RVT[19], DMANet[45], and HMNet[21], visualizations are built on top of their visualized image converted from event data directly by Metavision SDK of Prophesee because their latent feature maps contain much larger channels and smaller spatial resolutions, making them hard to visualize. For our TEDNet, visualizations are built on top of their latent feature maps $H^{t}$, which contain fewer channels and the same spatial resolutions as their input feature maps $I^{t}$. While the values scatter everywhere in latent feature maps $H^{t}$ due to D3D and hence visualization is difficult, we apply an adaptive histogram equalization method called CLAHE[54] to each channel in $H^{t}$ individually to enhance the contrast and choose one with the best appearance. For CenterTrack[50] and PermaTrack[43], visualizations are built on top of their input feature maps $I^{t}$, and CLAHE[54] is also applied to each channel in which one with the best appearance is chosen to be comparable with TEDNet.
@@
cell61 = c=4, rows = rowsep=0pt, columns = colsep=1pt, Ground Truth &
@@ 5 Conclusion
We introduce a novel architecture to apply tracking through occlusion to solve the long-range dependency problem for still objects in event cameras. The architecture retains the most abundant spatial-temporal information and tracks only still or pseudo-occluded objects instead of real occluded objects. Our comprehensive experiments demonstrate the effectiveness of each component of our TEDNet that achieves state-of-the-art performance in event-based object detection. In addition, we propose an auto-labeling algorithm to not only add additional labels for still objects to supervise the training of object permanence but also discard those impossible-to-detect objects at the very beginning of some videos to make model training more robust.
@@
In the supplementary material, we provide additional video and some qualitative analysis to demonstrate why our TEDNet outperforms prior works by such significant amounts.

### v3

@@ Abstract
Event-based object detection has recently garnered attention in the computer vision community due to the exceptional properties of event cameras, such as high dynamic range and no motion blur. However, feature asynchronism and sparsity cause invisible objects due to no relative motion to the camera [^1] The objects with no relative motion to the camera, still objects, invisible objects, and pseudo-occluded objects are used interchangeably in the paper., posing a significant challenge in the task. Prior works have studied various implicit-learned memories to retain as many temporal cues as possible. However, implicit memories still struggle to preserve long-term features effectively. In this paper, we consider those invisible objects as pseudo-occluded objects and aim to detect them by tracking through occlusions. Firstly, we introduce visibility attribute of objects and contribute an auto-labeling algorithm to not only clean the existing event camera dataset but also append additional visibility labels to it. Secondly, we exploit tracking strategies for pseudo-occluded objects to maintain their permanence and retain their bounding boxes, even when features have not been available for a very long time. These strategies can be treated as an explicit-learned memory guided by the tracking objective to record the displacements of objects across frames. Lastly, we propose a spatio-temporal feature aggregation module to enrich the latent features and a consistency loss to increase the robustness of the overall pipeline. We conduct comprehensive experiments to verify our method’s effectiveness where still objects are retained, but real occluded objects are discarded. The results demonstrate that (1) the additional visibility labels can assist in supervised training, and (2) our method outperforms state-of-the-art approaches with a significant improvement of 7.9% absolute mAP. Our code and the clean dataset with additional visibility labels are available at https://github.com/tkyen1110/TEDNet.
@@ 1 Introduction
Event cameras measure only per-pixel intensity changes asynchronously and are suitable for detecting fast-moving objects. For those slow-moving objects or still objects, it is very hard to notice their physical existence if there are no temporal cues. For example, Fig. 1 shows that the features in the ground truth bounding boxes disappear gradually when slowing down, become empty when being still, and reappear gradually when speeding up. Human beings are capable of knowing the physical existence of these two cars as long as they observe the whole video sequence, even if there are no features in the ground truth bounding boxes. This kind of ability is known as object permanence[3, 2, 10]. In fact, object permanence is a term from developmental psychology that describes how infants know the continuous existence of unseen objects.
@@
Prior works propose different kinds of memory architectures to retain object permanence in event-based object detection[34, 19, 26, 45, 21]. One common characteristic among these models is that they all use implicit memories with different spatial or temporal resolutions. After some experiments [^2] We cannot experiment on ASTMNet[26] due to no available source code. HMNet[21] is trained on GEN1 Automotive Detection Dataset with a smaller spatial resolution (304 x 240). Therefore, we downsample the spatial resolution of 1 Megapixel Automotive Detection Dataset (1280 x 720) and test them on HMNet[21] directly. shown in Fig. 1, we found that only RVT[19] can better handle the long-range dependency problem, i.e., objects become invisible when there are no relative motions between objects and the camera. Meanwhile, RVT[19] produces many false positives with high confidence scores due to its implicit ConvLSTMs. Other models in Fig. 1 are memoryless as long as the duration of missing features is too long, which is over 30 seconds from Fig. 1(c) to Fig. 1(d) and degrades the mAP performance of object detection.
@@
In order to supervise the training of object permanence, we need additional visibility labels to distinguish between moving objects and still objects. Hence, we propose an algorithm to perform the annotation automatically. Furthermore, our model is based on some joint object detection and tracking models, which give object detectors tracking ability. We augment these models with a spatio-temporal feature aggregation module and a consistency loss to not only enrich the latent features as shown in Fig. 1(g) but also increase the robustness of the overall pipeline where still objects are retained but real occluded objects are discarded. In summary, our model outperforms state-of-the-art event-based object detectors by 7.9% absolute mAP, and our main contributions are as follows:
@@
    We propose an innovative insight into memorizing missing features in event cameras by using tracking as an explicit memory instead of the traditional implicit memory method.
@@
    We propose an auto-labeling algorithm to provide a more reliable dataset for future works, leading to better training results without costly manual labor.
@@ 3 Method
Figure 2: The architecture of the proposed TEDNet. Spatio-Temporal Feature Aggregation integrates a 3D convolution with a recurrent neural network where $H_{t}=\emph{X3D}(I_{t},\ H_{t-1})$. Joint Object Detection and Tracking consists of localization head $f_{p}$/map $P_{t}$, size head $f_{s}$/map $S_{t}$, offset head $f_{off}$/map $O_{t}$, displacement head $f_{d}$/map $D_{t}$, visibility head $f_{v}$/map $V_{t}$, and consistency head $f_{c}$/map $C_{t}$. The novel consistency map $C_{t}$ correlates the proportional relationship between $D_{t}$ and $V_{t}$ where the novel consistency loss $L_{con}$ offers a better regularization to harness the advantage of X3D and increase the robustness of the overall pipeline.
@@ 3.1.3 Continuity Maintenance
is used to make the variation of visibility label for each specific tracking identity as smooth as possible and discard those bounding boxes with no features at their first occurrence. If there is no corresponding bounding box with the same tracking identity in the previous frame, a visibility label is determined by only the occupancy rate from Algorithm 1 to 1. Also, a buffer dubbed still_hits consists of several key-value pairs where keys are tracking identities, and values are initialized to 0 and limited to 5, representing the number of still hits for each tracking identity. still_hits can help to smooth out the impulse of visibility variation owing to the original label noises. If there is a corresponding tracked object in the previous frame, a visibility label is determined by either both occupancy rate and displacement from Algorithm 1 to 1 or only the still_hits buffer from Algorithm 1 to 1. Finally, not only visible objects but also those invisible objects that are tracked for two previous frames are added to the new labels from Algorithm 1 to 1 where track_vis consists of several key-value pairs where keys are tracking identities, and values are queues with size 2, representing the continuity of frame identities. As a result, only $N_{i}$ out of $M_{i}$ objects are preserved with additional visibility labels for $i$-th frame.
@@
Input: $inputFeatures$ = tensor of shape $[T,\ C,\ H,\ W]\ ;\ i\in[1,\ T]$
         $oldLabels[i]=\{[x_{j},\ y_{j},\ w_{j},\ h_{j},\ trackId_{j}]\ |\ j\in[1,\ M_{i}]\}$
@@
Output: $newLabels[i]=\{[x_{k},\ y_{k},\ w_{k},\ h_{k},\ trackId_{k},\ visibility_{k}]\ |\ k\in[1,\ N_{i}]\}$
@@ 26                   $still\_hits[trackId_{j}]\leftarrow still\_hits[trackId_{j}]+1$;
29            if $visibility_{j}=1.0$ or $track\_vis[trackId_{j}]=[i-2,\ i-1]$ then
@@ 3.3 Spatio-Temporal Feature Aggregation
The most significant challenge in event cameras is the long-range dependency problem. In this work, we propose to apply either a simple 3D convolutional network(C3D)[44] or a deformable 3D convolutional network(D3D)[49] as well as a ConvGRU in joint object detection and tracking module as our spatio-temporal feature aggregator. Unlike [44, 49] that input a video sequence to get a better spatio-temporal feature for the downstream task, our approach inputs only two frames, i.e., a current frame $I_{t}$ and a previous history frame $H_{t-1}$. The current history frame is obtained by $H_{t}=\emph{X3D}(I_{t},\ H_{t-1})$ where X3D is either C3D or D3D. To be more specific, we integrate an X3D with a recurrent neural network into our novel recurrent X3D to get a better spatio-temporal feature as an input to the following joint object detection and tracking module.
@@ 3.5 Consistency Loss
$min\Big{(}L_{con}\Big{)}=min\bigg{(}\Big{|}g(D_{t})\times h(C_{t})-V_{t}\Big{|}\bigg{)}=min\bigg{(}\Big{|}{\|D_{t}\|}_{2}\times e^{-relu(C_{t})}-V_{t}\Big{|}\bigg{)}$ (5)
$L_{con}=\frac{1}{N}\sum_{i=1}^{N}\bigg{|}{\Big{\|}D^{p_{i}}_{t}\Big{\|}}_{2}\times e^{-relu(C^{p_{i}}_{t})}-V^{p_{i}}_{t}\bigg{|}$ (6)
@@ 4.1 Dataset with Auto-Labeling
Prophesee, a company developing event-based sensors and algorithms, released the first large-scale and high-resolution (1280 $\times$ 720) event-based object detection dataset and its deep neural network RED[34] for automotive scenarios. This 1 Megapixel Automotive Detection Dataset is captured by the 1-megapixel event camera[17] and a standard RGB camera, but only event data and its annotations are released. The annotations come from the inference result of the RGB image and transfer to the event camera coordinate by geometric transformation. After visualizing the annotations in the dataset, we found that there are several incorrect bounding boxes with either imprecise locations and sizes or incorrect labels. Most importantly, several bounding boxes contain no features at the very beginning of some videos because they are still and visible by an RGB camera but invisible by an event camera, and humans cannot recognize their physical existence with only event data even though we have temporal cues. These incorrect bounding boxes limit the object detection performance and make model training awkward. Therefore, we clean the dataset by an auto-labeling algorithm in Sec. 3.1 with hyperparameters $D_{value}$=0.03 and $O_{value}$=0.1 to not only insert additional visibility labels to supervise the training of object permanence but also discard those bounding boxes that are impossible to be detected by humans. We named the dataset before/after data cleaning noisy GT/clean GT and its corresponding model noisy model/clean model.
@@ 4.2 Ablation Study
The models in ablation studies are trained on clean GT and tested on both noisy GT and clean GT as shown in Tab. 2. We choose two joint object detection and tracking models as our baselines, CenterTrack and PermaTrack[43], respectively.
@@
* The CenterTrack here is the original one[50] with ConvGRU, making it comparable
    with PermaTrack[43].
4.2.1 Spatio-Temporal Feature Aggregation
4.2.2 Consistency Loss
We conduct consistency loss between displacement map $D_{t}$ and visibility map $V_{t}$ to model their proportional relationship. Consistency loss can only be applied to PermaTrack[43] because there is no visibility head in CenterTrack[50]. Comparing (d) with (g), the consistency loss enhances the mAP@0.5 by 2.3% on noisy GT and 2.5% on clean GT. Comparing (f) with (h), the consistency loss degrades the mAP@0.5 by 0.9% on noisy GT and 0.4% on clean GT if we use C3D as the spatio-temporal feature aggregation module. In contrast, comparing (e) with our TEDNet, the consistency loss improves the mAP@0.5 by 2.4% on noisy GT and 2.3% on clean GT if we use D3D as the spatio-temporal feature aggregation module. To sum up, consistency loss offers a better regularization to harness the advantages of deformable convolutions and increases the robustness of the overall pipeline.
@@ 4.4 Visualization of Detection Results
We visualize the detection results in Fig. 3 by a video with two cars remaining still for over 30 seconds and some other cars passing in front of them in between. For RED[34], RVT[19], DMANet[45], and HMNet[21], visualizations are built on top of their visualized image converted from event data directly by Metavision SDK of Prophesee because their latent feature maps contain much larger channels and smaller spatial resolutions, making them hard to visualize. For our TEDNet, visualizations are built on top of their latent feature maps $H_{t}$, which contain fewer channels and the same spatial resolutions as their input feature maps $I_{t}$. While the values scatter everywhere in latent feature maps $H_{t}$ due to D3D and hence visualization is difficult, we apply an adaptive histogram equalization method called CLAHE[54] to each channel in $H_{t}$ individually to enhance the contrast and choose one with the best appearance. For CenterTrack[50] and PermaTrack[43], visualizations are built on top of their input feature maps $I_{t}$, and CLAHE[54] is also applied to each channel in which one with the best appearance is chosen to be comparable with TEDNet.
@@
cell62 = c=4, cell72 = c=4, rows = rowsep=0pt, columns = colsep=1pt, Ground Truth &
@@ 5 Conclusion
We introduce a novel architecture to apply tracking through occlusion to solve the long-range dependency problem for still objects in event cameras. The architecture retains the most abundant spatial-temporal information and tracks only still or pseudo-occluded objects instead of real occluded objects. Our comprehensive experiments demonstrate the effectiveness of each component of our TEDNet that achieves state-of-the-art performance in event-based object detection. In addition, the proposed auto-labeling algorithm provides a more reliable dataset and makes model training more robust.
Acknowledgements
This work was supported in part by National Science and Technology Council, Taiwan, under Grant NSTC 111-2634-F-002-022 and by Qualcomm through a Taiwan University Research Collaboration Project.
@@ Acknowledgements
In the supplementary material, we provide dataset details, implementation details, visualization of failure cases, additional video, and some quantitative analysis to demonstrate why our TEDNet outperforms prior works by such significant amounts.
1 Dataset Details
1 Megapixel Automotive Detection Dataset is the first large-scale and high-resolution (1280 $\times$ 720) event-based object detection dataset released by Prophesee, a company developing event-based sensors. The annotations of this dataset come from the inference result of the corresponding RGB image and are transferred to the event camera coordinate by geometric transformation, where only event data and its transformed annotations are released. The annotations consist of more than 25 million bounding boxes of cars, pedestrians, and two-wheelers recorded in 60Hz. In addition, bounding boxes of trucks, buses, traffic signs, and traffic lights are also provided. The dataset can be split into training, validation, and testing, with 11.19, 2.21, and 2.25 hours of recording, respectively. Each recording is split into 60-second chunks to increase the training and evaluation efficiency. To speed up the training, each annotation is preprocessed to record not only the bounding boxes at the current timestamp but also the corresponding bounding boxes at the previous timestamp as the ground truth of the tracking vector.
As mentioned in Sec. 3.1, an auto-labeling algorithm is proposed to not only clean the impossible-to-detect objects as shown in Fig. 4 but also add additional labels to distinguish between still and moving objects for the supervision of training object permanence as shown in Fig. 5.
{tblr}
row5 = font=, c, cell11 = c=5, cell31 = c=5, rows = rowsep=0pt, columns = colsep=1pt, Noisy Ground Truth (Noisy GT) &
Clean Ground Truth (Clean GT)
(a) 0.50 s (b) 17.00 s (c) 32.50 s (d) 33.25 s (e) 33.95 s
$\xrightarrow{\hskip 62.59596pt}$ time
Figure 4: A video sequence of the noisy and clean ground truth with absolute timestamps and ground truth labels. One car on the left remains still from the start of the video with no features, and human beings are unaware of the physical existence of this car. Hence, the bounding box of that car is removed according to the auto-labelling algorithm mentioned in Sec. 3.1. Bounding boxes with different colors correspond to different categories.
{tblr}
row3 = font=, c, cell11 = c=5, rows = rowsep=0pt, columns = colsep=1pt, Clean Ground Truth with Visibility Labels &
(a) 21.55 s (b) 21.80 s (c) 38.65 s (d) 55.50 s (e) 56.60 s
$\xrightarrow{\hskip 62.59596pt}$ time
Figure 5: A video sequence of the clean ground truth with absolute timestamps and visibility labels. Two cars slow down, and the visibility labels are changed from 1.0 (green) to 0.0 (red). Bounding boxes with different colors correspond to different visibility (mobility).
2 Implementation Details
Our TEDNet is built on top of PermaTrack where its original modules and hyperparameters are kept unchanged. Instead of training the model on sequences of length 17 with a batch size 16 in PermaTrack, TEDNet is trained on sequences of length 4 with a batch size 9. Each 60-second chunk is split into a sequence of length 1200, i.e., each time window contains 50 ms event data. The memory of our recurrent X3D and ConvGRU are reset at the end of each batch of chunks. The recurrent X3D has a feature dimension of 640 $\times$ 360 with 3 $\times$ 3 convolution kernels, and ConvGRU has a feature dimension of 160 $\times$ 90 with 7 $\times$ 7 convolution kernels. The learning rate scheduler is the same as the one used in PermaTrack where it is set to $1.25\times 10^{-4}$ and is decreased by a factor of 10 every 7 epochs for 1 epoch and increased back to the original value. All models are optimized using the Adam optimizer on 2 Nvidia 3090 GPUs with 100 epochs.
{tblr}
row15 = font=, c, cell11 = c=5, cell31 = c=5, cell51 = c=5, cell71 = c=5, cell92 = c=5, cell111 = c=5, cell131 = c=5, rows = rowsep=0pt, columns = colsep=1pt, Ground Truth &
RED 0.1
RED 0.4
PermaTrack 0.1
PermaTrack 0.4
TEDNet 0.1
TEDNet 0.4
(a) 53.15 s (b) 55.10 s (c) 55.35 s (d) 55.65 s (e) 56.85 s
$\xrightarrow{\hskip 62.59596pt}$ time
Figure 6: A video sequence of the failure case with absolute timestamps, ground truth labels, and some detection results with different confidence thresholds. One car on the left remains still for a while, and one motorcyclist passes between the still car and the camera. The values after model names are the thresholds where the bounding boxes with confidence scores lower than those thresholds are filtered out. Bounding boxes with different colors correspond to different categories.
3 Visualization of Failure Cases
We visualize the failure case in Fig. 6 by a video with one car on the left remaining still for a while and one motorcyclist passing between the still car and the camera. This scenario is different from the one in Sec. 4.4 where no intervention between the still objects and the camera. The root cause of the failure case in Fig. 6 comes from the degradation of confidence scores when the intervention occurs between the still objects and the camera. Meanwhile, the D3D in our TEDNet causes the ghosting effect of the motorcyclist in the 3rd and 4th frames of our TEDNet where the wrong and low-confidence-score bounding box is produced to detect the residual features. As a result, neither PermaTrack nor our TEDNet works well when the intervention occurs between the still objects and the camera. Namely, even though we regard those still and invisible objects as pseudo-occlusions, such that our TEDNet can track them and keep the bounding boxes, our TEDNet has weak capabilities to track real-occlusion with a high confidence score. However, the prior work RED can maintain the high-confidence-score bounding box even if the intervention occurs due to its implicit memory architecture and no feature missing for a very long time, which is different from Sec. 4.4.